OpenStudy (anonymous):

Sarah wants to hang a mirror in her room but the mirror and frame must not have an area larger than 15 square feet. The mirror is 3 feet wide and 4 feet long. Which quadratic equation represents the area of the mirror and frame combined? (The frame must have equal width of x on each side.)

OpenStudy (anonymous):

3x2 + 14x - 2 = 0 4x2 + 14x - 3 = 0 x2 + 12x - 14 = 0 2x2 + 7x - 14 = 0

OpenStudy (anonymous):

FLVS

OpenStudy (anonymous):

Well, here's the picture you should draw.